# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ 22 Lyapunov bound
$\mathbf{E}(Y^{r})=\sum_{i=0,\dots,r}{r\choose i}\frac{O(\frac{(n!)^{2i}(i!)^{2n}}{(ni)!})(\frac{n!}{m^{n}})^{r-i}(-1)^{r-i}}{m^{ni}}=\frac{1}{m^{nr}}\sum_{i=0,\dots,r}(-1)^{r-i}O(\frac{{r\choose i}(n!)^{r+i}(i!)^{2n}}{(ni)!}).$
Since $n\gg 1$, we can use a Stirling approximation for $i>0$
@@
$\mathbf{E}(Y^{r})=\frac{(\frac{n}{e})^{nr}}{m^{nr}}\bigg{(}(-1)^{r}O(1)(2\pi n)^{\frac{r}{2}}+\sum_{i=1,\dots,r}(-1)^{r-i}O\bigg{(}{r\choose i}(i!)^{2n}\frac{1}{i^{ni}}\frac{(2\pi n)^{\frac{i+r}{2}}}{\sqrt{2\pi ni}}\bigg{)}\bigg{)}.$
For any $i\leq r$, $\frac{(2\pi n)^{\frac{i+r}{2}}}{\sqrt{2\pi ni}}=O(n^{r})$, we can simplify the above to
$\mathbf{E}(Y^{r})=\frac{O(n^{r})(\frac{n}{e})^{nr}}{m^{nr}}\bigg{(}(-1)^{r}O(1)+\sum_{i=1,\dots,r}(-1)^{r-i}O\bigg{(}{r\choose i}(i!)^{2n}\frac{1}{i^{ni}}\bigg{)}\bigg{)}.$
Now, we will look at $|(-1)^{r}O(1)+\sum_{i=1,\dots,r}(-1)^{r-i}O({r\choose i}(i!)^{2n}\frac{1}{i^{ni}}))|=|(-1)^{r}O(1)+\sum_{i=1,\dots,r}(-1)^{r-i}O({r\choose i}(\frac{(i!)^{\frac{2}{i}}}{i})^{ni})|$. $\alpha(i):=\frac{(i!)^{\frac{2}{i}}}{i}$ is a monotonically increasing function of $i$ for $i>2$, this can be directly verified using Stirling’s approximation for large $i$, since $\frac{(i!)^{\frac{2}{i}}}{i}\approx\frac{O(i^{2})}{i}=O(i)$. For small $i$, we verified this by numerical simulation. Furthermore, $\alpha(i)\geq 1$ for $i\geq 1$. Therefore $|(-1)^{r}O(1)+\sum_{i=1,\dots,r}(-1)^{r-i}O({r\choose i}(\frac{(i!)^{\frac{2}{i}}}{i})^{ni})|=O((\alpha(r))^{nr})=O(\beta(r)^{n}),$ where $\beta(r):=\alpha(r)^{r}=\frac{(r!)^{2}}{r^{r}}.$ Note that $\beta(r)>1$ for $r>2$. Therefore
$|\mathbf{E}(Y^{r})|=\frac{O(n^{r}(\frac{n}{e})^{nr}\beta(r)^{n})}{m^{nr}}.$
@@
$\frac{N|\mathbf{E}(Y^{r})|}{\sigma^{r}_{N}}=O\bigg{(}\frac{\beta(r)^{n}}{N^{\frac{r}{2}-1}}\bigg{)}.$

### v3

@@ 16.1 Prospects of improving bounds on the interference terms
$\mathbf{E}(Y^{r})=\sum_{i=0,\dots,r}{r\choose i}\frac{A\frac{(n!)^{2i}(i!)^{2n}}{(ni)!}(\frac{n!}{m^{n}})^{r-i}(-1)^{r-i}}{m^{ni}}=\frac{1}{m^{nr}}\sum_{i=0,\dots,r}(-1)^{r-i}B_{i}\frac{{r\choose i}(n!)^{r+i}(i!)^{2n}}{(ni)!},$
for some constants $A,B_{i}>0$ for $i\in\{0,\dots,r\}$. Since $n\gg 1$, we can use a Stirling approximation for $i>0$
@@
$\mathbf{E}(Y^{r})=\frac{(\frac{n}{e})^{nr}}{m^{nr}}\bigg((-1)^{r}A(2\pi n)^{\frac{r}{2}}+\sum_{i=1,\dots,r}(-1)^{r-i}B_{i}{r\choose i}(i!)^{2n}\frac{1}{i^{ni}}\frac{(2\pi n)^{\frac{i+r}{2}}}{\sqrt{2\pi ni}}\bigg).$
For any $i\leq r$, $\frac{(2\pi n)^{\frac{i+r}{2}}}{\sqrt{2\pi ni}}\in O(n^{r})$, we can simplify the above to
$\mathbf{E}(Y^{r})=A\frac{n^{r}(\frac{n}{e})^{nr}}{m^{nr}}\bigg(B(-1)^{r}+\sum_{i=1,\dots,r}(-1)^{r-i}C_{i}{r\choose i}(i!)^{2n}\frac{1}{i^{ni}}\bigg),$
for some constants $A,B,C_{i}>0$. Now, we will look at $|B(-1)^{r}+\sum_{i=1,\dots,r}(-1)^{r-i}C_{i}{r\choose i}(i!)^{2n}\frac{1}{i^{ni}}))|=|(-1)^{r}B+\sum_{i=1,\dots,r}(-1)^{r-i}C_{i}{r\choose i}(\frac{(i!)^{\frac{2}{i}}}{i})^{ni}|$. $\alpha(i):=\frac{(i!)^{\frac{2}{i}}}{i}$ is a monotonically increasing function of $i$ for $i>2$, this can be directly verified using Stirling’s approximation for large $i$, since $\frac{(i!)^{\frac{2}{i}}}{i}\in\frac{O(i^{2})}{i}\in O(i)$. For small $i$, we verified this by numerical simulation. Furthermore, $\alpha(i)\geq 1$ for $i\geq 1$. Therefore $|(-1)^{r}B+\sum_{i=1,\dots,r}(-1)^{r-i}C_{i}{r\choose i}(\frac{(i!)^{\frac{2}{i}}}{i})^{ni}|\in O((\alpha(r))^{nr})\in O(\beta(r)^{n}),$ where $\beta(r):=\alpha(r)^{r}=\frac{(r!)^{2}}{r^{r}}.$ Note that $\beta(r)>1$ for $r>2$. Therefore
$|\mathbf{E}(Y^{r})|\in\frac{O(n^{r}(\frac{n}{e})^{nr}\beta(r)^{n})}{m^{nr}}.$
@@
$\frac{N|\mathbf{E}(Y^{r})|}{\sigma^{r}_{N}}\in O\bigg(\frac{\beta(r)^{n}}{N^{\frac{r}{2}-1}}\bigg).$